NOAA/NESDIS Banner
Office of Systems Development

Office of Satellite Operations Banner
Skip to contents    DOC   NOAA   NESDIS


Satellite Image of a storm forming.

 Find Past Report by Date

 


 Daily News Report

SATOPS Morning Report: Tuesday, January 28, 2014

Customers or users who want to call to ask about problems or report outages can use the following contacts:
- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-683-1400
- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 028

GOES
All GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
N15 rev#81693 Wallops: G2B did not lock up from spacecraft. Unable to retrieve data post-pass from Wallops. Phantom playback/data lost]

All other POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No other anomalies affecting product delivery or s/c health & safety.

SNPP
SNPP rev#11667 18:17 Marginal: TDRSS blind acquisition to test MX13 software.
At AOS White Sands DHN was not forwarding VCID1.
Ground system engineer recycled DHN and full real time telemetry was immediately restored.

All other Suomi NPP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more SNPP Information please see:
http://npp.gsfc.nasa.gov/

ESPC Reported data outage and delays
At DOY 028 0645z the GOES East and West images are not processing on LRIT. Issue is currently being investigated.

Questions or comments to Ron.Mahmot@noaa.gov

Last | Next | Previous | First


Home  |   Operations  |  News  |  Daily Updates |  Info & History
GOES |  POES  |  DMSP |  Contact Us   

Webmaster
Privacy Policy
NOAA link Dept. of Commerce link Link to OSO Homepage Link to NOAA Homepage